OpenStudy (anonymous):

Help: What are the intercepts of 3x − 9y = -27? (-3, 0) and (0, 9) (3, 0) and (0, 9) (3, 0) and (0, 6) (-9, 0) and (0, 3)

Theoceangirl (the_ocean_girl):

(-9,0) and (0,3)

OpenStudy (anonymous):

why do you pick that?

OpenStudy (onepieceftw):

Don't give answers, explain it to him.

OpenStudy (anonymous):

^

OpenStudy (onepieceftw):

A y intercept is the y value when x = 0 The x intercept is the x value when y = 0 plug in 0 for x and solve for y, that's the y intercept plug in 0 for y and solve for x, that's the x intercept

OpenStudy (anonymous):

Bruno what do you know so far?

OpenStudy (anonymous):

so its 3(0)-9(0)=-27?

OpenStudy (onepieceftw):

This is our equation: 3x − 9y = -27 Y intercept = 3(0) - 9y = -27 since anything ultiplied by 0 is 0 we have -9y = -27 -27/-9= y-intercept X intercept = 3x - 9(0) = -27 3x=-27 -27/3 = x-intercept

OpenStudy (anonymous):

ok I got it now thanks @OnePieceFTW

OpenStudy (onepieceftw):

No problem, bud